OpenStudy (anonymous):

Solve the following problem: Parking in the city lot can have at most 1/4 of the spaces allocated to compact cars. What is the least number of total spaces in the lot if there are 30 spaces allocated to compact cars? A. 200 B. 120 C. 180 D. 150

OpenStudy (callisto):

answer you need = 30 / (1/4) = work it out yourself :D

OpenStudy (callisto):

Or do it by an equation. Let x be the least number of total space x (1/4) = 30 x = 30/(1/4)